Staff:

LBTO: Jennifer (Luci), Doug, Juan Carlos (AO Support)

Observer: Enrico, et al (Arcetri remote)

Summary.

Forecast.

20230130_forecast0.png

20230129_forecasto0.png

Details.

Second half of the night for the DX Comm.

11:38PM telescope closed due to high wings.

LBTI handover to SoulDX earlier.

11:49PM

Jenny power on the WFS.

Telescope con stow pins in. authorize luci.

BCU47 (ccd47 not live).
  • REbin
  • OK.
OCAM2K not live.
  • Powercycle BCU2k+ rebin
  • OK

Google log.

LINK

-- JuanCarlosGuerra - 30 Jan 2023
Topic revision: r1 - 30 Jan 2023, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback